OpenStudy (anonymous):

A root of x 2 – 5x – 1 = 0 is

OpenStudy (whpalmer4):

That can't be factored into a product of two binomials. Do you know how to complete the square or use the quadratic formula?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thats a question on my accuplacer practice test i dont remember math that well

OpenStudy (whpalmer4):

Either of those would get you your two roots. Do you need a refresher on how to do them?

OpenStudy (whpalmer4):

I'll give you an example of both ways, but I'll do it with a different problem. Let's try \[x^2-4x-2 = 0\] Completing the square: To solve a quadratic by completing the square, we need to make sure that the coefficient of the x^2 term is 1. If it isn't, divide both sides of the equation by whatever it is to make it 1. Our x^2 has a coefficient of 1, so we don't need to do anything here. Write the equation with the x^2 and x terms on the left, and the number on the right. \[x^2 - 4x - 2 = 0\]Add 2 to both sides to move the 2 to the right side of the equals sign.\[x^2-4x -2 + 2 = 0 + 2\]\[x^2-4x = 2\] Next, take half of the coefficient of the x term, square it, and add it to both sides. \[x^2 - 4x + (\frac{-4}{2})^2 = 2 + (\frac{-4}{2})^2\] That's completing the square. Now we write the left side as a square: \[(x-(\frac{4}{2}))^2 = 2 + (\frac{16}{4})\]Which simplifies to \[(x-2)^2 = 6\]Now we take the square root of both sides which gives us\[(x-2) = \pm\sqrt{6}\]Solving for \(x\) \[x = 2 \pm \sqrt{6}\]We have to use the \(\pm\) in front of the square root because there are two solutions to a quadratic. If you plug each of the solutions we got into the equation and expand it, you'll see that both of them give a result of 0. The quadratic formula: The solution to \(ax^2+bx+c=0\) is given by\[x = \frac{-b\pm\sqrt{b^2-4ac}}{2a}\]With our sample problem, we have a = 1, b = -4, c = -2. Plugging them into the formula gives us\[x=\frac{-(-4)\pm\sqrt{(-4)^2-4(1)(-2)}}{2(1)} = \frac{4\pm\sqrt{16-(-8)}}{2} = 2\pm\frac{\sqrt{24}}{2}\]We can factor a 4 out of the square root\[x=2\pm\frac{\sqrt{4*6}}{2} = 2\pm\frac{2\sqrt{6}}{2} = 2\pm\sqrt{6}\] Just like we got by completing the square.

OpenStudy (whpalmer4):

Now you have a set of tools to solve any quadratic you might encounter.

OpenStudy (anonymous):

thank you now do u have an email if i get stuck i really need a tutor and no one I know is smarter than me